# User Simulation for Evaluating Information Access Systems

### v1

@@ 8.3 Extending Current User Simulation Technologies for Evaluation
User simulation for recommender systems. Even though we intended to cover both search engines and recommender systems in the book, most of the content in the previous chapters has been conducted in the context of search engines. We took the stance that search engines and recommender systems are “two sides of the same coin” and there are indeed several shared elements, e.g., when a user interacts with a ranked list of recommended information items the similarities. This is also reflected in the fact IR evaluation methodology has been increasingly adopted for evaluating recommender systems. However, there is undoubtedly a much better understanding of user behaviour when it comes to interacting with search engines than in the case of recommender systems, which is an obvious future opportunity to address.
@@
User simulation as MDP policy learning. Another limitation of existing work is that it focuses on developing and evaluating specific techniques, while there is very little work on developing general formal models for user simulation. Earlier we attempted to use an MDP as a general formal framework to bring together separate lines of research in user simulation. There is clearly a need for doing more research along this direction. Specifically, future research may use MDPs as a formal framework in which specific simulators may be instantiated by learning a policy for each state of the MDP framework.
Infrastructures for user simulation-based evaluation. Evaluation using user simulation requires infrastructures that go beyond traditional test collections. As we briefly discussed in Sections 6.7 and 7.5, the current work in this area is still quite limited, which may also be a reason why user simulation has not yet been widely used for evaluation. Thus, there is a need for further research and development of scalable and sustainable infrastructures for supporting user simulation-based evaluation. This includes, e.g., releasing toolkits with implementations of user simulators, setting up web-based service systems that researchers can use directly to evaluate their interactive systems, systematically archived experiment results on user simulators, and open-source infrastructures to enable massive collaboration of researchers in contributing new user simulators. Such infrastructures would also naturally facilitate collaboration between industry and academia, which is required for sustaining the innovation and delivering broader impact in this area.
@@ 8.4 Long-term Challenges and Future Research
8.4.1 Realisticity
The first general challenge lies in a more rigorous definition of the problem of user simulation itself. While informally we can easily understand what it means by simulating a user computationally, which naturally implies that an ideal user simulator should be as similar to a real user as possible, how to rigorously and mathematically define the problem remains a major open challenge. In particular, the question “How to make simulators more realistic?” needs to be formally specified as some kind of optimization problem. Viewed in the MDP framework, this general challenge has to do with the overall formulation of the problem, i.e., how to define an agent with all the necessary components in the MDP, including the definition of state, actions, reward, and the policy to be simulated.
The importance of this problem was already recognized by the community as one of the open questions identified at the Sim4IR workshop [Balog:2022:SIGIRForum] concerned the realisticity of simulators. The following quote from the workshop report explains well this important foundational question about user simulation:
@@ 8.4.1 Realisticity
This touches on the issue of formulating the problem of user simulation and evaluating user simulation itself. Depending on the purpose and the anticipated use of a user simulator, we may want to frame and evaluate the problem of user simulation differently. There are inevitably many trade-offs that have to be made, such as interpretability vs. prediction accuracy, task-specific simulation of users vs. general simulation, individual user simulation vs. user group simulation. Practical factors, such as data availability and the possibility of overfitting models to specific observed users in a dataset, are also important to consider.
@@
For example, for the purpose of data augmentation, the user simulators do not have to be interpretable and thus do not have to be built based on realistic user models; instead, a user simulator would be acceptable if the generated data are realistic. This, however, raises the question about how do we quantify the similarity of the generated data with real user data? Is it acceptable as long as it is very close to one user’s data or does it have to be close to the data from a group of users? For analyzing user behaviour, a realistic user simulator would have to be based on realistic user models. In such a case, interpretability is important, whereas how well the generated data match empirical data observed from a real user may be secondary. This of course raises another open question which is how to rigorously determine whether a model is realistic.
@@
The trade-off between interpretability and prediction accuracy is reflected in the wide spectrum of possible computational models and algorithms for user simulation already developed, which included both highly interpretable, but overly simplistic models (e.g., the cascade model for scanning and clicking) and purely data-driven models that can fit empirical user rating/clicking behaviour well (e.g., deep neural networks for predicting clicks in recommender systems), with many models in between. A major variable here is the nature of parameters used in the model. When all the parameters are interpretable, we would have not just one single user simulator but a whole family of user simulators that can vary meaningfully to model individual differences of users in the family by varying some of the parameter values. However, due to the constraints on the interpretability of parameters, the model’s predictive power would also be inevitably limited (by the parameters allowed), thus their empirical performance in predicting user behaviour would generally be not as high as the pure data-driven models, where the empirical accuracy of prediction is an explicitly defined objective function to be optimized. Ideally, in the future, we should integrate these two kinds of models by using increasingly more interpretable parameters, which may be viewed as deeper understanding of the users as well, since such interpretable user simulators can be regarded as theoretical claims about users. The data-driven models could be used to understand the gap between a theoretical model and empirical observations, i.e., reveal what kind of patterns the theoretical models fail to explain. Those can then become a research problem to be further studied, leading to new research results that can eventually help bridge the gap with additional menaningful parameters introduced to the simulator or additional relations captured between those latent user variables.
@@ 8.4.2 Evolving System Functions and User Interfaces
The second general challenge is that a full specification of a user simulator must include a full characterization of the system functions and interfaces that a user would interact with. However, both the functions and user interfaces vary across systems and they also evolve over time. For example, while early-generation search systems used simple list-based interfaces, modern SERPs and recommendations are organized into slates, i.e., non-linear layouts. There is also the question of conversational UIs, which support multiple modalities. Since the actions available for a user to take generally depend on the interfaces and functions of a system, this variation creates a major challenge in developing general user simulation techniques, since the most effective simulation techniques for a user action may depend on the whole user interface or other functions that the user had access to.
@@
A particularly interesting general question here is to what extent we can develop general policies that might be invariant to the variations of user interfaces. This is presumably feasible for simulating some user actions, since there may be some general psychological mechanism that all/most users use. For example, despite many variations of how a list of search result would be shown exactly to a user, from a user’s perspective, perhaps a sequential browsing (linear traversing) model would often be used by the user to prioritize the scanning of results; the difference would be more reflected in the clicking decisions, which in turn might be affected more by how exactly a snippet is shown. The challenge here is how to decompose a complex policy in user simulation into multiple potentially generalizable component policies. In some sense, the current work has already made such a decomposition when query simulation and clicking simulation are studied separately. However, it is unclear who such a decomposition is to be performed in principled way and how common components and workflows from multiple systems are to be extracted. One possibility may be to to analyze the interaction data collected from multiple systems and use data mining approaches, such as frequent pattern mining, or latent variable models, such as hidden Markov models, to discover empirically meaningful common interface components. This way, we may create some basic user simulators designed based on some common interfaces or interface components shared by many systems, and use them as a prior for adaptation to a specific system for construction of a more system-specific simulator. A major challenge here is determining what kind of parameters would need to be introduced to enable customization and how to set the values of those parameters.
@@ 8.4.3 Modeling Variable Latent States of Users: Cognitive and Emotional States
The third general challenge lies in the modeling of latent states of users, including, e.g., cognitive states and emotional states. Since the observable user behaviours are often a result of a latent state of the user, modeling such latent states are crucial for accurate user simulation. However, by definition, such latent states are generally not observable (directly), which makes their modeling a difficult challenge. While there has been some work on modeling users’ knowledge and cognitive state, and simulating query formulation and clicking based on these, the current work has often used very simple models (e.g., unigram language models). Also, the knowledge state is generally limited to a user’s familiarity with relevant and non-relevant content. In the future, more research is needed in this area.
@@
First, for the simulation of any user interacting with a information access system, there is a need to model the user’s information need. The ASK theory [Belkin:1982:JD] provides some basis for modeling the case when the user has a real need for some information that they are unfamiliar with; it can be used to conveniently build a comprehensive model of information need by leveraging knowledge graphs that are widely available nowadays. According to ASK, an information need reflects an anomalous state of knowledge (ASK), and thus we might assume that part of a knowledge graph is “unavailable" to a user and the user’s goal (task) is to obtain relevant information to the missing part, thus filling the “hole” in the user’s knowledge graph. However, in a broader context of search applications, such as product search, we also often have users search for known items (e.g., purchasing a consumed product regularly). In that scenario, ASK might not be the right framework as the user might have excellent knowledge about relevant content. There, the information need is more about identifying the exact information item(s) in a collection that matches well with what the user has already had in mind. (We might still consider there is some “hole" in the user’s knowledge graph, which, e.g., is the exact URL to the product that the user wanted to purchase.) In other cases, a user’s information need may be quite complex (e.g., in medical search), and the user may also face a significant vocabulary gap. Then, modeling how a user learns to bridge the vocabulary gap during the search process would be required in order to accurately simulate the user’s behaviour. In general, for a complex information need, we will need complex variables to characterize a user’s information need accurately, along with ways to capture the associated (and inevitable) uncertainty. Moreover, a user’s information need likely evolves during the search process as the user is being exposed to more information items, making it necessary to model how the information need would be updated over time.
@@
Second, a user’s search behaviour depends much on the user’s knowledge background in general. Thus, another essential element in any user simulator is a model of the user’s background knowledge. Of particular importance is to model the user’s familiarity with relevant and non-relevant content. This would allow user simulation to be employed for both expert and novice users, as well as any users in between. To make such a background knowledge model meaningful, there should be some additional interpretable categorical variables associated with users. For example, elementary students might be modeled using a similar background knowledge model, which would clearly be different from that used for graduate student users. To make the background knowledge models realistic, such models can be estimated based on the real content produced by a certain group of users. In the extreme case, if we can collect content generated by an individual user, we may use all the observed content from the user to build a background model for the user. As in the case of an information need, a user’s knowledge would also evolve over time during the search process, from which the user can be expected to gain new knowledge. This requires modeling the cognitive process of a user during search and updating the user’s knowledge model based on the new information the user has been exposed to. An interesting connection may be made here between user simulation and work on “search as learning” [Vakkari:2016:JIS].
@@
Third, a user’s search behaviour depends on the user’s past experiences with the system. For example, if the system tends to perform better with keyword queries as opposed to natural language questions, then users will learn to issue the former. Another important component of a user simulator, therefore, is a model of a user’s knowledge about how the system works. In some sense, the PRE framework for query formulation is based on the assumption that a user would imagine what would happen to a candidate query if that query were to be executed by the system, and then try to assess the expected Recall and Precision of the candidate query (i.e., assessing the optimality). However, there is no explicit model about how a system works proposed in the work so far. In the future, modeling and updating a user’s knowledge about the search system would be necessary. How to develop such models and incorporate them into user simulation is a highly interesting challenge for future research.
@@ 8.4.4 Evaluation of User Simulators: Validity and Reliability
As already touched upon in the discussion of realisticity earlier, how to evaluate user simulators is yet another major difficult challenge. In theory, utilizing user simulation for evaluation seems logical; however, in practice, the reliability of this evaluation methodology is called into question if the user simulators themselves lack accuracy. Given that the current measures are using essentially quite naive user simulators, one might assume that imperfect user simulators would also be acceptable. However, the real challenge here is to quantify the reliability of a user simulator, so that we at least would know the limitations when we use a certain kind of user simulators for evaluation. Compared with the work on developing user simulation techniques, there has been little research on how to evaluate the user simulators themselves. The Tester-based framework [Labhishetty:2021:SIGIR] provides a basis for evaluating user simulators using a set of necessary requirements, but those requirements are indeed far from sufficient. The reliability estimation algorithm proposed in [Labhishetty:2022:ECIR] would allow for assigning reliability scores to a set of Testers and user simulators simultaneously. This may provide us with an initial step toward quantifying the validity and reliability of user simulators. However, how to develop more Testers is also a new challenge that must be tackled before we can have a sufficiently large set of Testers for evaluating user simulators.
@@
To ensure realistic user modeling, there is no way to escape from the measure of similarity between a user simulator and an actual user. Here we may distinguish different perspectives of similarity. First, we might consider model similarity. In such a case, we would have to have user models available for us to compare again. Realistically, we may only have available a component model for a specific action or behaviour of a user, though as the current understanding of a user is still quite limited. With the availability of any component user model, we can then compare the model used in a simulator with the component user model. Second, we might consider data similarity. In such a case, we would compare the simulator-generated data with the data observed from real users. The comparison can be done from different perspectives. A natural decomposition to compute the overall similarity score based on aggregation over similarity scores on each action to be simulated; see our discussion earlier under Realisticity (Section 8.4.1.)
@@
In sum, there are many challenges associated with the evaluation of user simulators. While human involvement for validate is inevitable, the question of how is might be minimized is an important one for obvious reasons.
8.5 User Simulation as an Interdisciplinary Research Field
AI systems can be broadly categorized into two types: autonomous AI, where machines aim to perform tasks normally performed by humans (e.g., self-driving cars), and assistive AI, where the goal is not to replace humans, but rather aid them to fulfill their tasks more efficiently and effectively, often by augmenting human intelligence. A main challenge unique to assistive AI systems is how to collaborate with a human user effectively so as to maximize the combined intelligence of humans and machines. Today, some of the most successful instances of assistive AI are search engines, recommender systems, and conversational assistants. The book has focused on such special cases of assistive AI systems, but most of discussions are also applicable to assistive AI systems in general, thus opening up many interesting future research and collaboration opportunities spanning multiple research communities.
From another perspective, user simulation can be used not only for evaluation but also for other applications as we briefly discussed earlier in the book. Different applications may require or appreciate different kinds of simulation techniques, but there is clear synergy of those different lines of exploration. In the future, it is desirable to integrate those scattered efforts to advance the general technologies for simulating users, which again requires interdisciplinary research involving multiple research communities.
We thus envision that the major challenges we discussed above will likely attract researchers from different research areas, making the topic of user simulation a promising interdisciplinary area for research intersecting with many related subject areas, including information science, information retrieval, recommender systems, machine learning, natural language processing, knowledge representation, human-computer interaction, and psychology.
8.5.1 User Simulation and Machine Learning
As an application where machine learning (ML) can be applied, user simulation provides many interesting opportunities for applying a broad variety of ML techniques, including unsupervised learning, supervised learning, imitation learning, and reinforcement learning. Generative probabilistic models may be applied to learn patterns of user actions in an unsupervised way. The observed user actions can be used as labels for use with supervised learning or imitation learning to build a user simulator. Reinforcement learning is necessary for user simulation since it enables a user simulator to optimize sequential decision making (policy) by learning from its interactions with a system. User simulation can also be used to generate synthetic user (interaction) data which can be useful for training machine learning algorithms, especially complex algorithms, such as reinforcement learning, that often requires either massive amounts of user interaction data or an interactive environment for training. User simulation techniques would enable the training of such algorithms that are currently infeasible.
All such knowledge also needs to be updated to reflect learning by a search system user during the search process [Vakkari:2016:JIS]. Therefore, it is necessary to study the cognitive processes of searchers. This makes the study of user simulation closely related to intelligent tutoring systems, where simulated learners have also been used for evaluating intelligent education systems [Dorcca:2015:AIED]. In general, how to represent the knowledge that a user possesses, how to model the reasoning a user makes over the knowledge, and how to capture a user’s knowledge gain during a search process are all important yet challenging open questions that may require deeper understanding of a fundamental question in AI—how knowledge should be represented formally.
8.5.3 User Simulation and HCI and Psychology
Psychology and HCI research play a vital role in establishing the theoretical foundations for simulating users, making them not only closely linked to user simulation but also essential for constructing realistic simulators. Looking ahead, we anticipate a growing integration of knowledge and theories from psychology and HCI into the development of user simulators. Leveraging any insights gained about users has the potential to enhance the accuracy and fidelity of user simulation.
8.5.4 User Simulation and Specific User Tasks
Domain-specific or system-specific user simulation would require collaborative research between user simulation researchers and domain experts. The incorporation of domain-specific knowledge and assumptions would be necessary. Since information access systems are widely used by many people in all kinds of application domains, there will be widespread opportunities for collaboration with domain experts, leading to many specialized user simulators developed for specific domains. For example, in the case of simulating searches in the medical domain, we likely would need to incorporate specific knowledge about medical search tasks, whereas in e-commmerce search, it would be necessary to incorporate other kinds of domain knowledge, depending on the specific type of products to be purchased by the user.
8.5.5 User Simulation and Software Systems
Finally, in contrast to the static test collections currently used for evaluation, user simulators need to function as operational software platforms to facilitate evaluation. Consequently, an interesting relationship emerges between user simulation and software systems, as the challenge lies in constructing “high-performance” user simulators capable of simultaneously interacting with multiple search systems (to enable their evaluation). Given that a sophisticated user simulator will likely have to use advanced algorithms and data structures (e.g., for representation of knowledge) and support large-scale experiments (that possibly explore various configurations), there will be significant new challenges in improving the efficiency of user simulators.

### v2

@@ 10.3 Extending Current User Simulation Technologies for Evaluation
User simulation for recommender systems. Even though we intended to cover both search engines and recommender systems in the book, most of the content in the previous chapters has been conducted in the context of search engines; user simulation for the evaluation of recommender systems has only been studied recently [Rahdari:2022:IntRS, Hazrati:2024:UMUAI, Ghanem:2022:ECRA, Rahdari:2024:TORS]. We took the stance that search engines and recommender systems are “two sides of the same coin” and there are indeed several shared elements, e.g., when a user interacts with a ranked list of recommended information items the similarities. This is also reflected in the fact IR evaluation methodology has been adopted and extended for evaluating recommender systems [Cremonesi:2010:RecSys, Parapar:2021:RecSys]. However, there is undoubtedly a much better understanding of user behaviour when it comes to interacting with search engines than in the case of recommender systems, which is an obvious future opportunity to address.
@@
Modeling user simulation as an MDP. Another limitation of existing work is that it focuses on developing and evaluating specific techniques, while there is very little work on developing general formal models for user simulation. Earlier we attempted to use an MDP as a general formal framework to bring together separate lines of research in user simulation. There is clearly a need for doing more research along this direction. Specifically, future research may use MDPs as a formal framework in which specific simulators may be instantiated by specifying states, actions, rewards, and policy.
Infrastructures for user simulation-based evaluation. Evaluation using user simulation requires infrastructures that go beyond traditional test collections. As we briefly discussed in Section 8.4, the current work in this area is still quite limited, which may also be a reason why user simulation has not yet been widely used for evaluation. Thus, there is a need for further research and development of scalable and sustainable infrastructures for supporting user simulation-based evaluation. This includes, for example, releasing toolkits with implementations of user simulators, establishing web-based services that researchers can readily use to evaluate their interactive systems, maintaining a repository of experimental results on user simulators, and developing open-source infrastructures to facilitate large-scale collaboration among researchers in contributing new user simulators. Such infrastructures would also naturally facilitate collaboration between industry and academia, which is required for sustaining the innovation and delivering broader impact in this area.
@@ 10.4 Long-Term Challenges and Future Research
10.4.1 Realism
The first general challenge lies in a more rigorous definition of the problem of user simulation itself. While informally we can easily understand what it means by simulating a user computationally, which naturally implies that an ideal user simulator should be as similar to a real user as possible, how to rigorously and mathematically define the problem remains a major open challenge. In particular, the question “How to make simulators more realistic?” needs to be formally specified as some kind of optimization problem. Viewed in the MDP framework, this general challenge has to do with the overall formulation of the problem, i.e., how to define an agent with all the necessary components in the MDP, including the definition of states, actions, rewards, and the policy to be simulated.
The importance of this problem was already recognized by the community as one of the open questions identified at the Sim4IR workshop [Balog:2022:SIGIRForum] concerned the realism of simulators. The following quote from the workshop report explains well this important foundational question about user simulation:
@@ 10.4.1 Realism
This touches on the issue of formulating the problem of user simulation and evaluating user simulation itself. Depending on the purpose and the anticipated use of a user simulator, we may want to frame and evaluate the problem of user simulation differently. There are inevitably many trade-offs that have to be made, such as interpretability vs. prediction accuracy, task-specific simulation of users vs. general simulation, individual user simulation vs. user group simulation [Nguyen:2019:UMUAI]. Practical factors, such as data availability and the possibility of overfitting models to specific observed users in a dataset, are also important to consider.
@@
For example, for the purpose of data augmentation, the user simulators do not have to be interpretable and thus do not have to be built based on realistic user models; instead, a user simulator would be acceptable if the generated data are realistic. This, however, raises the question about how we should quantify the similarity of the generated data with real user data. Is it acceptable as long as it is very close to one user’s data or does it have to be close to the data from a group of users? For analyzing user behaviour, a realistic user simulator would have to be based on realistic user models. In such a case, interpretability is important, whereas how well the generated data match empirical data observed from a real user may be secondary. This of course raises another open question, which is how to rigorously determine whether a model is realistic.
@@
The trade-off between interpretability and prediction accuracy is reflected in the wide spectrum of possible computational models and algorithms for user simulation already developed, which included both highly interpretable, but overly simplistic models (e.g., the cascade model for scanning and clicking) and purely data-driven models that can fit empirical user rating/clicking behaviour well (e.g., deep neural networks for predicting clicks in recommender systems), with many models in between. A major variable here is the nature of parameters used in the model. When all the parameters are interpretable, we would have not just one single user simulator but a whole family of user simulators that can vary meaningfully to model individual differences of users in the family by varying some of the parameter values. However, due to the constraints on the interpretability of parameters, the model’s predictive power would also be inevitably limited (by the parameters allowed), thus their empirical performance in predicting user behaviour would generally be not as high as that of pure data-driven models, where the empirical accuracy of prediction is an explicitly defined objective function to be optimized. Ideally, in the future, we should integrate these two kinds of models by using increasingly more interpretable parameters, which may be viewed as deeper understanding of the users as well, since such interpretable user simulators can be regarded as theoretical claims about users. The data-driven models could be used to understand the gap between a theoretical model and empirical observations, i.e., reveal what kind of patterns the theoretical models fail to explain. Those can then become a research problem to be further studied, leading to new research results that can eventually help bridge the gap with additional menaningful parameters introduced to the simulator or additional relations captured between those latent user variables.
@@ 10.4.2 Evolving System Functions and User Interfaces
The second general challenge is that a full specification of a user simulator must include a full characterization of the system functions and interfaces that a user would interact with. However, both the functions and user interfaces vary across systems and they also evolve over time. For example, while early-generation search systems used simple list-based interfaces, modern recommender system organize items into slates and SERPs have non-linear layouts. There is also the question of conversational UIs, which support multiple modalities. Since the actions available for a user to take generally depend on the interfaces and functions of a system, this variation creates a major challenge in developing general user simulation techniques, since the most effective simulation techniques for a user action may depend on the whole user interface or other functions that the user has access to.
@@
A particularly interesting general question here is to what extent we can develop general policies that might be invariant to the variations of user interfaces. This is presumably feasible for simulating some user actions, since there may be some general psychological mechanism that all/most users follow. For example, despite many variations of how a list of search results would be shown exactly to a user, from a user’s perspective, perhaps a sequential browsing (linear traversing) model would often be used by the user to prioritize the scanning of results; the difference would be more reflected in the clicking decisions, which in turn might be affected more by how exactly a snippet is shown. The challenge here is how to decompose a complex policy in user simulation into multiple potentially generalizable component policies. In some sense, the current work has already made such a decomposition when query simulation and clicking simulation are studied separately. However, it is unclear how such a decomposition is to be performed in principled way and how common components and workflows from multiple systems are to be extracted. One possibility may be to analyze the interaction data collected from multiple systems and use data mining approaches, such as frequent pattern mining, or latent variable models, such as hidden Markov models, to discover empirically meaningful common interface components. This way, we may create some basic user simulators designed based on some common interfaces or interface components shared by many systems, and use them as a prior for adaptation to a specific system for construction of a more system-specific simulator. A major challenge here is determining what kind of parameters would need to be introduced to enable customization and how to set the values of those parameters.
@@ 10.4.3 Modeling Variable Latent States of Users: Cognitive and Emotional States
The third general challenge lies in the modeling of latent states of users, including, e.g., cognitive states and emotional states. A series of studies by Kuhlthau have examined information retrieval from both cognitive and affective aspects [Kuhlthau:1988:RQ, Kuhlthau:1991:JASIST]. To model a user holistically, it is important to model both their cognitive and emotional states, as these are interconnected. Both states can be latent, making them challenging to model directly. However, observable user behaviours often result from (or are determined by) specific latent states. Thus, modeling these latent states is crucial for accurately predicting a user’s observable behaviour.
@@
While there has been some work on modeling users’ knowledge and cognitive state, and simulating query formulation and clicking based on these, the current work has often used very simple models (e.g., unigram language models). Also, the knowledge state is generally limited to a user’s familiarity with relevant and non-relevant content. In the future, more research is needed in this area.
@@
First, for the simulation of any user interacting with a information access system, there is a need to model the user’s information need. The ASK theory [Belkin:1982:JD] provides some basis for modeling the case when the user has a real need for some information that they are unfamiliar with; it can be used to conveniently build a comprehensive model of information need by leveraging knowledge graphs that are widely available nowadays. According to ASK, an information need reflects an anomalous state of knowledge (ASK), and thus we might assume that part of a knowledge graph is “unavailable” to a user and the user’s goal (task) is to obtain relevant information to the missing part, thus filling the “hole” in the user’s knowledge graph. However, in a broader context of search applications, such as product search, we also often have users search for known items (e.g., purchasing a consumed product regularly). In that scenario, ASK might not be the right framework as the user might have excellent knowledge about relevant content. There, the information need is more about identifying the exact information item(s) in a collection that matches well with what the user has already had in mind. (We might still consider there is some “hole” in the user’s knowledge graph, which, e.g., is the exact URL to the product that the user wanted to purchase.) In other cases, a user’s information need may be quite complex (e.g., in medical search), and the user may also face a significant vocabulary gap. Then, modeling how a user learns to bridge the vocabulary gap during the search process would be required in order to accurately simulate the user’s behaviour. In general, for a complex information need, we will need complex variables to characterize a user’s information need accurately, along with ways to capture the associated (and inevitable) uncertainty. Moreover, a user’s information need likely evolves during the search process as the user is being exposed to more information items, making it necessary to model how the information need would be updated over time.
@@
Second, a user’s search behaviour depends much on the user’s knowledge background in general. Thus, another essential element in any user simulator is a model of the user’s background knowledge. Of particular importance is to model the user’s familiarity with relevant and non-relevant content. This would allow user simulation to be employed for both expert and novice users, as well as any users in between. To make such a background knowledge model meaningful, there should be some additional interpretable categorical variables associated with users. For example, elementary school students might be modeled using a similar background knowledge model, which would clearly be different from that used for graduate student users. To make the background knowledge models realistic, such models can be estimated based on the real content produced by a certain group of users. In the extreme case, if we can collect content generated by an individual user, we may use all the observed content from the user to build a background model for them. As in the case of an information need, a user’s knowledge would also evolve over time during the search process, from which the user can be expected to gain new knowledge. This requires modeling the cognitive process of a user during search and updating the user’s knowledge model based on the new information the user has been exposed to. An interesting connection may be made here between user simulation and work on “search as learning” [Vakkari:2016:JIS, Collins:2017:Report].
@@
Third, a user’s search behaviour depends on the user’s past experiences with the system. For example, if the system tends to perform better with keyword queries as opposed to natural language questions, then users will learn to issue the former. Another important component of a user simulator, therefore, is a model of a user’s knowledge about how the system works. In some sense, the PRE framework for query formulation [Labhishetty:2022:ICTIR] is based on the assumption that a user would imagine what would happen to a candidate query if that query were to be executed by the system, and then try to assess the expected Recall and Precision of the candidate query (i.e., assessing the optimality). However, there is no explicit model about how a system works proposed in the PRE framework so far. In the future, modeling and updating a user’s knowledge about the search system would be necessary. How to develop such models and incorporate them into user simulation is a highly interesting challenge for future research.
@@ 10.4.4 Evaluation of User Simulators: Validity and Reliability
As already touched upon in the discussion of realism earlier, how to evaluate user simulators is yet another major difficult challenge. In theory, utilizing user simulation for evaluation seems logical; however, in practice, the reliability of this evaluation methodology is called into question if the user simulators themselves lack accuracy. Given that the current measures are using essentially quite naive user simulators, one might assume that imperfect user simulators would also be acceptable. However, the real challenge here is to quantify the reliability of a user simulator, so that we at least would know the limitations when we use a certain kind of user simulators for evaluation. Compared with the work on developing user simulation techniques, there has been little research on how to evaluate the user simulators themselves. The Tester-based framework [Labhishetty:2021:SIGIR] provides a basis for evaluating user simulators using a set of necessary requirements, but those requirements are indeed far from sufficient. The reliability estimation algorithm proposed in [Labhishetty:2022:ECIR] would allow for assigning reliability scores to a set of Testers and user simulators simultaneously. This may provide us with an initial step toward quantifying the validity and reliability of user simulators. However, how to develop more Testers is also a new challenge that must be tackled before we can have a sufficiently large set of Testers for evaluating user simulators.
@@
To ensure realistic user modeling, there is no way to escape from the measure of similarity between a user simulator and an actual user. Here we may distinguish different perspectives of similarity. First, we might consider model similarity. In such a case, we would have to have user models available for us to compare again. Realistically, we may only have available a component model for a specific action or behaviour of a user though, as the current understanding of users is still quite limited. With the availability of any component user model, we can then compare the model used in a simulator with the component user model. Second, we might consider data similarity. In such a case, we would compare the simulator-generated data with the data observed from real users. The comparison can be done from different perspectives. A natural decomposition to compute the overall similarity score based on aggregation over similarity scores on each action to be simulated; see our discussion earlier under Realism (Section 10.4.1.)
@@
In sum, there are many challenges associated with the evaluation of user simulators. While human involvement for validation is inevitable, the question of how it might be minimized is an important one for obvious reasons.
Acknowledgements
We are grateful to the editors-in-chief, Yiqun Liu and, especially, Pablo Castells, for handling our manuscript. We are also extremely grateful to our two anonymous reviewers: their valuable comments and suggestions helped us improve the book.
The authors thank Nolwenn Bernard for her invaluable assistance in setting up and maintaining the annotated bibliography on the companion website. The authors further want to thank Craig Macdonald, Maria Vlachou, To Eun Kim, Ryen White, Yaxiong Wu, and Yong Li for their suggestions on literature to include.
Krisztian Balog was partially supported by the Norwegian Research Center for AI Innovation, NorwAI (Research Council of Norway, project number 309834). ChengXiang Zhai was partially supported by the US National Science Foundation and the Institute of Education Sciences through Award # 2229612 (National AI Institute for Inclusive Intelligent Technologies for Education).